# Scaling Laws for Floating–Point Quantization Training

### v1

@@ Abstract
Low-precision training is considered an effective strategy for reducing both training and downstream inference costs. Previous scaling laws for precision mainly focus on integer quantization, which pay less attention to the constituents in floating-point quantization and thus cannot well fit the LLM losses in this scenario. In contrast, while floating-point quantization training is more commonly implemented in production, the research on it has been relatively superficial. In this paper, we thoroughly explore the effects of floating-point quantization targets, exponent bits, mantissa bits, and the calculation granularity of the scaling factor in floating-point quantization training performance of LLM models. While presenting an accurate floating-point quantization unified scaling law, we also provide valuable suggestions for the community: (1) Exponent bits contribute slightly more to the model performance than mantissa bits. We provide the optimal exponent-mantissa bit ratio for different bit numbers, which is available for future reference by hardware manufacturers; (2) We discover the formation of the critical data size in low-precision LLM training. Too much training data exceeding the critical data size will inversely bring in degradation of LLM performance; (3) The optimal floating-point quantization precision is directly proportional to the computational power, but within a wide computational power range, we estimate that the best cost-performance precision lies between 4-8 bits.
Machine Learning, ICML
[…]
Equal contribution
@@ 1 Introduction
The various scaling laws of large language models (LLMs) could help developers effectively select superior parameter settings before experiments and accurately predict the model performance under different configurations, which are regarded as excellent guidance in LLM training. The widely acknowledged scaling law efforts such as Kaplan et al. (2020), Hoffmann et al. (2022), and Li et al. (2024) mainly concentrate on the central factors, i.e., model size and trained token size, which significantly impact the performance of LLMs. With the rapid growth of both model and data sizes, there has been increasing attention to the efficiency and cost of LLM training. Training and serving with lower precision becomes a popular solution. Currently, lots of representative LLMs are trained in BF16 and even lower precision (Dubey et al., 2024; Sun et al., 2024; Liu et al., 2024; Yang et al., 2024; Ma et al., 2024; Wang et al., 2023; Peng et al., 2023), aiming to balance effectiveness and efficiency. Compared to integer quantization, floating-point (FP) quantization can better maintain LLMs’ accuracy at extremely lower bit rates and thus is often equipped in low-precision LLMs. Therefore, exploring the scaling laws of LLM performance under different low precision settings with floating-point quantization becomes essential to shed light on future low-precision LLM training.
@@
Recently, there is pioneer work that conducts in-depth analyses and explorations on the LLM’s scaling laws for precision in both training and inference (Kumar et al., 2024), quantitatively measuring the degradation rules of post-train quantization and quantized training. This scaling law provides an appropriate conclusion explaining the potential damage of excessively increasing training data to low-precision LLMs’ performance. However, Kumar et al. (2024) directly adopts the bit width as the precision in its low-precision scaling laws, which may lose finer-grained modeling of the relationship between various parameter settings related to the FP quantization and the final loss of LLMs. In practice, the key factors of FP quantization such as the exponent, mantissa, and the block size of scaling factors may have different impacts on the final loss. A more comprehensive, precise, and practical scaling law for float quantized training related to the data size (D), model size (N), exponent (E), mantissa (M), and block size of scaling factors (B) is urgently desired.
@@
Our work concentrates on establishing, verifying, and analyzing the scaling law for float quantized training in LLMs. At the beginning, we first predict the model performance via the precision-related scaling law from previous work under different data/model sizes and precision settings. Surprisingly, we discover that the predictive performance was not perfectly satisfactory under different float quantized training settings. Subsequently, we carefully design a comprehensive set of explorations with experiments of different precision settings (training 366 models), exploring the basic scaling law formation, as well as the potential impact of the quantization targets, exponent and mantissa, and block sizes on the loss. Finally, we aggregate these factors to get our final scaling law for float quantized training with valuable insights to guide the LLM training under low precision. Our scaling law is formulated as follows:
@@
The first two factors $D$ and $N$ indicate the data size and model size respectively, which show the main impacts on training loss given by the key factors of data and model size similar with the Chinchilla scaling law (Hoffmann et al., 2022). $\epsilon$ represents the bias. The last factor could be regarded as the additional negative impact deriving from low precision training, where $\frac{D^{\beta}}{N^{\alpha}}$ implies a certain form of “knowledge intensity” in LLM, and $\log_{2}B$, $(E+0.5)^{\delta}$, and $(M+0.5)^{\nu}$ jointly reflect the “low precision information loss” of float quantized training. We have conducted extensive fitting experiments with various possible scaling law formulations to ensure the accuracy and simplicity of our scaling laws. Note that the exponential hyper-parameters $\alpha$ and $\beta$ of model and data sizes are exactly the same as those in the first two factors. The product of the above “knowledge intensity” and “low precision information loss” forms the last factor.
Figure 1 and Figure 8 illustrates the fitting results of our scaling law compared with other existing scaling laws, demonstrating our advantages on predicting LLM performances under different float quantized training settings. Throughout our experiments and analyses related to our scaling law, we also discover the following observations and insights that could facilitate future low-precision LLM training: (a) It has been discovered that the impact of quantized weights on the performance is relatively minor during both forward and backward computations. Meanwhile, activations demonstrate a higher degree of quantization tolerance specifically when computing gradients pertaining to themselves. (b) The data size of LLM pre-training cannot be added indefinitely without harming the performance under low precision, while large model sizes, higher precision settings (measured by exponent and mantissa), and smaller block sizes could increase the extreme point of effective trained tokens for LLM training. (c) Intuitively, the negative impact of low-precision training in LLMs will be proportionally amplified with the “knowledge intensity”. (d) The exponent and mantissa have their optimal settings under different bit widths. Exponent bits contribute slightly more to the model performance than mantissa bits. (e) The optimal floating-point quantization precision exhibits a direct proportionality with computational power. Nonetheless, across a broad spectrum of computational power, our estimated optimal cost-performance precision resides within the 4-8 bits range.
@@ 2 Preliminary
Classical Scaling Laws. Scaling laws have become a fundamental framework for understanding the relationship between essential factors such as model size (N), data size (D), and the resulting loss (L) in deep learning. Two classical scaling laws have been widely recognized in the industry: the Chinchilla scaling law (Hoffmann et al., 2022) and the OpenAI scaling law (Kaplan et al., 2020). The Chinchilla scaling law is expressed as:
@@
On the other hand, the OpenAI scaling law is given by:
@@
Scaling Laws for Precision. Subsequent research extends this framework by incorporating the role of precision in quantized training and inference, providing insights into how precision affects model performance. In Kumar et al. (2024), the precision-aware scaling laws are introduced, capturing the trade-offs between model size $N$, data size $D$, and precision $P$. For integer quantized training, they propose the tradeoff between weight $N$ and weight precision $P$ as:
@@
Current Scaling Laws cannot Well Fit in Floating-point Quantization. While most prior work has focused on integer quantized training, floating-point quantization is more prevalent in real-world applications due to its hardware compatibility and finer granularity. For instance, formats such as FP16 and BF16 are standard in many large-scale training pipelines, and emerging formats like FP8 and FP4 are gaining traction. Despite this, scaling laws specifically tailored to floating-point quantization are still largely unexplored. The primary distinction between floating-point and integer quantization lies in the allocation and usage of bits. Floating-point numbers allocate bits to represent both the exponent and the mantissa, with each set of bits serving distinct purposes: the exponent mainly captures dynamic range, while the mantissa mainly encodes precision within that range. In contrast, integer formats uniformly distribute all bits to refine the quantization lattice, providing consistent resolution across the representable range. This fundamental difference highlights the need for dedicated scaling laws for the unique characteristics of floating-point formats.
Figure 1: The fitting results of the scaling law in Eq. (7) deriving from Kumar et al. (2024), which have large bias in E1M1 case. In the three sub-figures on the left, middle and right, the sizes of the data points are approximately proportional to $D$, $E$, and $M$ respectively.
Kumar et al. (2024) hypothesizes that the exponent and mantissa bits are scaled jointly (i.e., increase together as total bit count does). Then, in floating-point formats, precision is determined by the exponent $E$ and mantissa $M$, with the total precision:
$P=E+M+1.$ (6)
By substituting $P$ in their precision-aware scaling law, we have:
@@
However, upon conducting experiments and applying this scaling law to fit empirical results, particularly for low-precision training regimes, we observed significant deviations between the law’s predictions and actual performance, as illustrated in Figure 1. The unsatisfactory fit, especially for training results using low-bit floating-point formats, suggests that the previous proposed relationship in Kumar et al. (2024) does not adequately capture the nuanced dynamic impacts of floating-point quantization on LLM performance.
In this work, we address these shortcomings by re-deriving the scaling law for floating-point quantized training. Our re-derivation incorporates a more nuanced understanding of how the finer factors of exponent, mantissa, and block size affect low-precision training. By refining the theoretical framework and aligning it more closely with observed behaviors, we aim to establish a more accurate and predictive scaling law that bridges the gap between theoretical insights and real-world applications.
@@ 3.1 Method and Implementation
3.1.1 Quantization Method
We quantize a tensor into a low-precision floating-point format, following the IEEE 754 standard (Kahan, 1996), which includes both normal and subnormal representations. The format consists of a sign bit, $E$ exponent bits and $M$ mantissa bits. To expand the dynamic range, the special bits are adopted for normal values instead of representing Infinity and Not a Number (NaN). Since the modern hardware does not support arbitrary floating-point format, we simulate them using QPyTorch (Zhang et al., 2019) with nearest rounding.
Due to the narrow dynamic range and low representation precision of the low-precision format, we employ scaling techniques (Sun et al., 2019; Micikevicius et al., 2022). The original tensor is multiplied by a higher-precision scaler before being cast into the low-precision format. The scaling factor is computed as follows:
$S_{i}=\frac{\text{FP}_{\text{max}}}{\text{max}\left(\lvert\mathbf{X_{\left[Bi:B\left(i+1\right)\right]}}\rvert\right)},$ (8)
where ${\text{FP}_{\text{max}}}$ represents the maximum normal value of the low-precision floating-point format. A scaling factor can be shared every $B$ elements along the channel dimension. It is a unified representation for tensor-wise scaling ($B=bd_{\text{in}}$), channel-wise scaling ($B=d_{\text{in}}$) and block-wise ($1\leq B\textless d_{\text{in}}$) scaling for a tensor with the shape of $b\times d_{\text{in}}$.
3.1.2 Implementation
The quantization is applied to the linear layers in transformer (Vaswani, 2017), excluding the dot-product attention and the classifier. The computation of a linear layer involves a matrix multiplication (Eq. (9)) during the forward phase and two matrix multiplications (Eq. (10) and Eq. (11)) during the backward phase.
$\mathbf{Y}=\mathbf{X}\mathbf{W}^{T},$ (9)
$\mathrm{d}\mathbf{X}={\mathrm{d}\mathbf{Y}_{1}}{\mathbf{W}_{\text{bwd}}},$ (10)
$\mathrm{d}\mathbf{W}={\left(\mathrm{d}\mathbf{Y}_{2}\right)}^{\mathrm{T}}\mathbf{X}_{\text{bwd}},$ (11)

### v3

@@ Abstract
Low-precision training is considered an effective strategy for reducing both training and downstream inference costs. Previous scaling laws for precision mainly focus on integer quantization, which pay less attention to the constituents in floating-point (FP) quantization, and thus cannot well fit the LLM losses in this scenario. In contrast, while FP quantization training is more commonly implemented in production, it’s research has been relatively superficial. In this paper, we thoroughly explore the effects of FP quantization targets, exponent bits, mantissa bits, and the calculation granularity of the scaling factor in FP quantization training performance of LLM models. In addition to an accurate FP quantization unified scaling law, we also provide valuable suggestions for the community: (1) Exponent bits contribute slightly more to the model performance than mantissa bits. We provide the optimal exponent-mantissa bit ratio for different bit numbers, which is available for future reference by hardware manufacturers; (2) We discover the formation of the critical data size in low-precision LLM training. Too much training data exceeding the critical data size will inversely bring in degradation of LLM performance; (3) The optimal FP quantization precision is directly proportional to the computational power, but within a wide computational power range. We estimate that the best cost-performance precision should lie between 4-8 bits.
@@ 1 Introduction
Scaling laws of large language models (LLMs) could help developers effectively select superior parameter settings before experiments and accurately predict the model performance under different configurations. They are regarded as excellent guidance in LLM training. The widely acknowledged scaling law efforts such as Kaplan et al. (2020), Hoffmann et al. (2022), and Li et al. (2024) mainly concentrated on the central factors, i.e., model size and trained token size, which significantly impact the performance of LLMs. With the rapid growth of both model and data sizes, there has been increasing attention to the efficiency and cost of LLM training. Training and serving with lower precision becomes a popular solution. Currently, lots of representative LLMs were trained in BF16 and even lower precision (Dubey et al., 2024; Sun et al., 2024; Liu et al., 2024; Yang et al., 2024; Ma et al., 2024; Wang et al., 2023; Peng et al., 2023), aiming to balance effectiveness and efficiency. Compared to integer quantization, floating-point (FP) quantization can better maintain LLMs’ accuracy at extremely lower bit rates and thus is often equipped in low-precision LLMs. Therefore, exploring the scaling laws of LLM performance under different low precision settings with FP quantization becomes essential to shed light on future low-precision LLM training.
@@
Recently, there was pioneer work that conducted in-depth analyses and explorations on the LLM’s scaling laws for precision in both training and inference (Kumar et al., 2024), quantitatively measuring the degradation rules of post-train quantization and quantized training. This scaling law provides an appropriate conclusion explaining the potential damage of excessively increasing training data to low-precision LLMs’ performance. However, Kumar et al. (2024) directly adopted the bit width as the precision in its low-precision scaling laws, which might lose finer-grained modeling of the relationship between various parameter settings related to the FP quantization and the final loss of LLMs. In practice, the key factors of FP quantization such as the exponent, mantissa, and the block size of scaling factors may have different impacts on the final loss. A more comprehensive, precise, and practical scaling law for FP quantized training related to the data size (D), model size (N), exponent (E), mantissa (M), and block size of scaling factors (B) is urgently desired.
@@
Our work concentrates on establishing, verifying, and analyzing the scaling law for FP quantized training in LLMs. At the beginning, we first predict the model performance via the precision-related scaling law from previous work under different data/model sizes and precision settings. Surprisingly, we discover that the predictive performance was not perfectly satisfactory under different FP quantized training settings. Subsequently, we carefully design a comprehensive set of explorations with experiments of different precision settings (training 366 models), exploring the basic scaling law formation, as well as the potential impact of the quantization targets, exponent and mantissa, and block sizes on the loss. Finally, we aggregate these factors to get our final scaling law for FP quantized training with valuable insights to guide the LLM training under low precision.
Our FP quantization training scaling law, namely Capybara (Appendix B), is formulated as follows:
@@
The first two factors $D$ and $N$ indicate the data size and model size respectively, which show the main impacts on training loss given by the key factors of data and model size similar to the Chinchilla scaling law (Hoffmann et al., 2022); $\epsilon$ represents the bias; The last factor could be regarded as the additional negative impact deriving from low precision training, where $\frac{D^{\beta}}{N^{\alpha}}$ implies a certain form of “knowledge intensity” in LLM, and $\log_{2}B$, $(E+0.5)^{\delta}$, and $(M+0.5)^{\nu}$ jointly reflect the “low precision information loss” of FP quantized training. We have conducted extensive fitting experiments with various possible scaling law formulations to ensure the accuracy and simplicity of our scaling laws. Note that the exponential hyper-parameters $\alpha$ and $\beta$ of model and data sizes are exactly the same as those in the first two factors. The product of the above “knowledge intensity” and “low precision information loss” forms the last factor.
(a) Kumar et al. (2024).
(b) Ours.
Figure 1: Comparing Eq. (6) from Kumar et al. (2024) and Eq. (13) from our work, our Capybara scaling law fits data better in low - precision scenarios. Specifically, Kumar et al. (2024)’s fitting results show considerable bias in the E1M1 case. In the subfigures, data point magnitudes are roughly proportional to $E$.
Figure 1 illustrates the fitting results of our Capybara scaling law compared with other ones, demonstrating our advantages on predicting LLM performances under different float quantized training settings. Throughout our experiments and analyses related to our Capybara scaling law, we also discover the following observations and insights that could facilitate future low-precision LLM training: (a) It has been discovered that the impact of quantized weights on the performance is relatively minor during both forward and backward computations. Meanwhile, activations demonstrate a higher degree of quantization tolerance specifically when computing gradients pertaining to themselves. (b) The data size of LLM pre-training cannot be added indefinitely without harming the performance under low precision, while large model sizes, higher precision settings (measured by exponent and mantissa), and smaller block sizes could increase the extreme point of effective trained tokens for LLM training. (c) Intuitively, the negative impact of low-precision training in LLMs would be proportionally amplified with the “knowledge intensity”. (d) The exponent and mantissa have their optimal settings under different bit widths. Exponent bits contribute slightly more to the model performance than mantissa bits. (e) The optimal FP quantization precision exhibits a direct proportionality with computational power. Nonetheless, across a broad spectrum of computational power, our estimated optimal cost-performance precision should reside within the 4-8 bits range.
@@ 2 Preliminary
Classical Scaling Laws. Scaling laws have become a fundamental framework for understanding the relationship between essential factors such as model size (N), data size (D), and the resulting loss (L) in deep learning. Two classical scaling laws have been widely recognized in the industry: Chinchilla scaling law (Hoffmann et al., 2022) and OpenAI scaling law (Kaplan et al., 2020). The Chinchilla scaling law is expressed as:
@@
The OpenAI scaling law is given by:
@@
Scaling Laws for Precision. Subsequent research extends this framework by incorporating the role of precision in quantized training and inference, so as to provide insights into how precision affects model performance. In Kumar et al. (2024), precision-aware scaling laws were introduced to capture the trade-offs between model size $N$, data size $D$, and precision $P$. For integer quantized training, they proposed the tradeoff between weight $N$ and weight precision $P$ as:
@@
Current Scaling Laws cannot Fit Well in FP Quantization. Note that most previous work focused on integer quantized training. FP quantization is more prevalent in real-world applications due to its hardware compatibility and finer granularity. For instance, formats such as FP16 and BF16 are standard in many large-scale training pipelines, and emerging formats like FP8 and FP4 are gaining traction. Despite this, scaling laws specifically tailored to FP quantization are still largely unexplored. The primary distinction between FP and integer quantization lies in the allocation and usage of bits. FP numbers allocate bits to represent both the exponent and the mantissa, with each set of bits serving distinct purposes: the exponent mainly captures dynamic range, while the mantissa mainly encodes precision within that range. In contrast, integer formats uniformly distribute all bits to refine the quantization lattice, providing consistent resolution across the representable range. This fundamental difference highlights the need for dedicated scaling laws for the unique characteristics of FP formats.
Kumar et al. (2024) hypothesized that the exponent and mantissa bits should be scaled jointly (i.e., increase together as total bit count does). Then, in FP formats, precision is determined by the exponent $E$ and mantissa $M$, with the total precision: $P=E+M+1$. By substituting $P$ in their precision-aware scaling law, we have:
@@
However, upon conducting experiments and applying this scaling law to fit empirical results, particularly for low-precision training regimes, we observed significant deviations between the law’s predictions and actual performance, as illustrated in Figure 1(a). The unsatisfactory fit, especially for training results using low-bit FP formats, suggests that the previous relationship proposed in Kumar et al. (2024) does not adequately capture the nuanced dynamic impacts of FP quantization on LLM performance.
In this work, we address these shortcomings by re-deriving the scaling law for FP quantized training. Our re-derivation incorporates a more nuanced understanding of how the finer factors of exponent, mantissa, and block size affect low-precision training. By refining the theoretical framework and aligning it more closely with observed behaviors, we aim to establish a more accurate and predictive scaling law so as to bridge the gap between theoretical insights and real-world applications.
@@ 3.1 Method and Implementation
Quantization Method. We quantized a tensor into a low-precision FP format, following the IEEE 754 standard (Kahan, 1996), which includes both normal and subnormal representations. The format consists of a sign bit, $E$ exponent bits and $M$ mantissa bits. To expand the dynamic range, the special bits are adopted for normal values instead of representing Infinity and Not a Number (NaN). Since the modern hardware does not support arbitrary FP format, we simulate them using QPyTorch (Zhang et al., 2019) with nearest rounding. Due to the narrow dynamic range and low representation precision of the low-precision format, we employ scaling techniques (Sun et al., 2019; Micikevicius et al., 2022). The original tensor is multiplied by a higher-precision scaler before being cast into the low-precision format. The scaling factor is computed as follows:
$S_{i}=\text{FP}_{\text{max}}/\text{max}\left(\lvert\mathbf{X_{\left[Bi:B\left(i+1\right)\right]}}\rvert\right),$ (7)
where ${\text{FP}_{\text{max}}}$ represents the maximum normal value of the low-precision FP format. A scaling factor can be shared every $B$ elements along the channel dimension. It is a unified representation for tensor-wise scaling ($B=bd_{\text{in}}$), channel-wise scaling ($B=d_{\text{in}}$) and block-wise ($1\leq B\textless d_{\text{in}}$) scaling for a tensor with the shape of $b\times d_{\text{in}}$.
Implementation. The quantization is applied to the linear layers in transformer (Vaswani et al., 2017), excluding the dot-product attention and the classifier. In a linear layer’s computation, there is one matrix multiplication during the forward phase and two during the backward phase as:
$\begin{cases}\mathbf{Y}&=\mathbf{X}\mathbf{W}^{T}\\ \mathrm{d}\mathbf{X}&={\mathrm{d}\mathbf{Y}_{1}}{\mathbf{W}_{\text{bwd}}}\\ \mathrm{d}\mathbf{W}&={\left(\mathrm{d}\mathbf{Y}_{2}\right)}^{\mathrm{T}}\mathbf{X}_{\text{bwd}},\end{cases}$ (8)
@@
Modeling Object. As shown in Eq. (37) of (Kuzmin et al., 2022) , using the Signal-to-Quantization-Noise Ratio (SQNR) for GEMM operations could unify the effects of exponent (E) and mantissa (M) precision while accounting for input distribution modifications like random orthogonal transformations. However, SQNR calculation inherently depends on tensor distributions, which may limit its practical applicability. Therefore, we directly select the raw exponent and mantissa bits rather than the SNR as essential factors in our Capybara scaling law for more precise prediction ability.